### v1

# Transfer Learning for Paediatric Sleep Apnoea Detection Using Physiology-Guided Acoustic Models

Abstract
Paediatric obstructive sleep apnoea (OSA) is clinically significant yet difficult to diagnose, as children poorly tolerate sensor-based polysomnography. Acoustic monitoring provides a non-invasive alternative for home-based OSA screening, but limited paediatric data hinders the development of robust deep learning approaches. This paper proposes a transfer learning framework that adapts acoustic models pretrained on adult sleep data to paediatric OSA detection, incorporating SpO2-based desaturation patterns to enhance model training. Using a large adult sleep dataset (157 nights) and a smaller paediatric dataset (15 nights), we systematically evaluate (i) single- versus multi-task learning, (ii) encoder freezing versus full fine-turning, and (iii) the impact of delaying SpO2 labels to better align them with the acoustics and capture physiologically meaningful features. Results show that fine-tuning with SpO2 integration consistently improves paediatric OSA detection compared with baseline models without adaptation. These findings demonstrate the feasibility of transfer learning for home-based OSA screening in children and offers its potential clinical value for early diagnosis.
Index Terms—  Sleep Disordered Breathing, Paediatric Obstructive Sleep Apnoea, Acoustic Model, Transfer Learning, Physiological Delay Modelling
1 Introduction
Obstructive sleep apnoea (OSA) is a common paediatric sleep-disordered breathing condition characterised by repeated airway obstruction, leading to intermittent hypoxia, disrupted ventilation, and fragmented sleep [1]. Its prevalence is 1–5% in the general paediatric population [2], rising to 33–61% among children with obesity [3]. Untreated OSA is associated with cardiovascular, metabolic, and neurocognitive impairments, behavioural problems, growth delay, and reduced quality of life [4].
Fig. 1: Pipeline for data collection and model development. Breathing sounds are recorded using a smartphone, optionally supplemented with chest sensor and IMU signals. Clinician-annotated HST provides apnoea–hypopnoea events and AHI scores. Solid lines denote data used in the current model, while dashed lines indicate signals reserved for future integration. The fine-tuning pipeline loads an adult-pretrained multi-task CNN, optionally freezes the shared CNN backbone, and fine-tunes task-specific layers.
Polysomnography (PSG) is the diagnostic gold standard, but requires overnight lab-based monitoring with specialised equipment and personnel, resulting in high costs and low tolerability in children. While treatments are well established [5, 6], PSG complexity limits its utility as a universal diagnostic tool, especially for habitually snoring children at risk of OSA [7]. Consequently, over 90% of OSA cases in children remain undiagnosed [8], highlighting the need for accessible, affordable, minimally invasive diagnostics [9]. Alternative approaches, such as questionnaires [10], symptom-based scoring [11], and simplified recordings [12], lack sufficient diagnostic accuracy [13]. Machine learning (ML) shows promise for automated, objective paediatric OSA diagnosis [14], and acoustic monitoring offers a low-cost, non-invasive option for home-based OSA screening [15, 16]. However, paediatric applications remain limited due to weaker respiratory sounds, developmental physiology variability, and scarcity of labelled datasets.
OSA data collection is constrained by the slow and labour-intensive process of manual scoring. Paediatric data are particularly scarce, as children have lower tolerance for sensors, which further limits model development. Transfer learning [17] provides a practical solution by adapting models pretrained on large adult datasets to the paediatric domain. While successfully applied to sleep stage prediction from adults to children [18], its use in acoustic-based OSA detection and integration of physiological information, such as oxygen desaturation, has not yet been explored.
This study proposes transfer learning for paediatric sleep apnoea detection by adapting a convolutional neural network (CNN) trained on adult acoustic data. The framework supports both a single-task head predicting OSA and a multi-task head predicting OSA and time below a night-specific oxygen saturation (SpO2) threshold. Two adaptation strategies are investigated: freezing the encoder versus full fine-tuning. Physiological priors are incorporated by modelling the temporal delay between nasal airflow cessation and oxygen desaturation. Our contributions are threefold: (i) a transfer learning framework for acoustic-based paediatric OSA detection, (ii) evaluation of single- versus multi-task learning with different fine-tuning strategies with integration of physiological priors, and (iii) demonstration that transfer learning enables scalable, non-invasive OSA screening using smartphone-based acoustics without wearable sensors.
The rest of the paper is organised as follows. Section 2 describes the data used in this study. Section 3 presents the proposed transfer learning framework for acoustic-based paediatric OSA detection and integration with SpO2 information. Experiments are described in Section 4. Results and Discussion are given in Section 5. Section 6 concludes this paper and presents future directions.
2 Data
Paediatric sleep data were collected by Sheffield Children’s NHS Foundation Trust with ethics approval from the UK NHS Research Ethics Committees (REC reference: 24/PR/0284). 15 children (ages 1–15 years, mean $6\pm 4$; 5 females, 10 males; mean BMI $25\pm 10$) were recruited for overnight monitoring using a Home Sleep Apnoea Test (HSAT) device (SOMNOtouch™ RESP [^1] https://somnomedics.de/en/solutions/sleep_diagnostics/polygraphy-devices/somnotouch-resp/) and concurrent smartphone audio recordings in both laboratory and home settings (participant demographics in Table 1).
The HSAT measured SpO2, heart rate, airflow, respiratory effort, snoring, sleep–wake status, and body position, with an optional thorax sensor. Audio was recorded at 16 kHz/16-bit resolution in 2-minute segments via a custom mobile app (Figure 1). Apnoea events were manually scored by a registered polysomnographic technologist [^2] https://www.brpt.org/. In total, $\sim$120 hours of data were obtained across 15 nights. Audio quality was assessed by the snore-to-non-snore ratio, while HSAT signal quality was evaluated by the percentage of missing SpO2.
The adult dataset comprised 103 participants (67 males, 36 females; mean age 45 $\pm$ 13 years; mean BMI 31 $\pm$ 7), providing 157 nights and $\sim$1094 hours of recording (see [15] for more details).
3 Methods
Fig. 2: Time delay of oxygen desaturation and Mel-frequency spectrum. Red dashed rectangles mark the 15 s oxygen desaturation window, which is shifted from 0 to longer time delays to identify the delay that best discriminates between normal and abnormal sleep.
Fig. 3: Violin plots showing time spent below SpO2 baseline as a percentage within a 15 s window versus time delay of SpO2 for non-OSA, OSA and hypopnoea segments.
3.1 Feature extraction and preprocessing
Full-night audio recordings are segmented into overlapping 30 s windows $\mathcal{S}_{i}$ with a 10 s shift. Each segment is processed via short-term Fourier transform (50 ms Hann window, 20 ms hop) to obtain a power spectrogram $P(t,f)$. A 64-filter Mel-filterbank is applied to produce log-compressed Mel-spectrograms $M(t,m)$, followed by normalisation per Mel-bin to obtain $\hat{M}_{i}\in\mathbb{R}^{T\times F}$. These features serve as input to a CNN $f_{\theta}$ for binary OSA detection $y_{i}\in\{0,1\}$, with $y_{i}=1$ segments grouped into events and aligned with HSAT annotations. For physiological integration, oxygen desaturation is quantified as the percentage of time below the night-specific SpO2 baseline (3% drop) according to AASM guideline [^3] American Academy of Sleep Medicine guideline: https://aasm.org/clinical-resources/practice-standards/practice-guidelines/ within a 15 s window for each 30 s segment (see Fig. 2).
3.2 Problem formulation
Let the adult dataset be
$\mathcal{D}_{A}=\{(X_{i}^{A},y_{i}^{A},s_{i}^{A})\}_{i=1}^{N_{A}},$ (1)
where $X_{i}^{A}\in\mathbb{R}^{T\times F}$ is a Mel spectrogram segment, $y_{i}^{A}\in\{0,1\}$ is the OSA label, and $s_{i}^{A}\in[0,1]$ is the night-specific SpO2 time percentage, $N_{A}$ is number of night recordings. Let the paediatric dataset be
$\mathcal{D}_{C}=\{(X_{j}^{C},y_{j}^{C},s_{j}^{C})\}_{j=1}^{N_{C}},\quad N_{C}\ll N_{A}.$ (2)
[…]
$\Delta t=t_{\text{OOD}}-t_{\text{ONAC}},$ (3)
where $t_{\text{ONAC}}$ is the onset of nasal airflow cessation and $t_{\text{OOD}}$ is the onset of a 3% oxygen desaturation.
This work transfers an adult-trained model to the paediatric domain by systematically evaluating the impact of different strategies, including task type, fine-tuning approach, and the use of SpO2 labels. Research questions are formulated as:
RQ1: Does fine-tuning, particularly in multi-task learning, improve paediatric OSA detection compared to single-task learning and models without fine-tuning?
RQ2: How do fine-tuning strategies (frozen encoder, full fine-tuning) and the use of SpO2 labels (delayed versus immediate labels) influence model performance in paediatric OSA detection?
RQ3: Which combination of task type, fine-tuning strategy and SpO2 labels provides the most reliable and clinically meaningful predictions for paediatric OSA severity?
3.3 Physiological information
Analysis of paired acoustic and SpO2 segments revealed that OSA events show larger desaturation (below 3% oxygen desaturation) proportions than non-OSA events, and that incorporating a temporal delay enhances discrimination (see Fig. 3). Consistent with prior physiological findings [19], a median time delay of 26 s was observed across patients (range: 0–36 s). This delay was embedded into the multi-task learning pipeline, with both global and subject-specific delays evaluated to assess their impact on classification performance.
3.4 Transfer learning strategy
We decompose the adult CNN into an encoder $g_{\theta}$ that extracts feature embeddings from a spectrogram segment, and a prediction head $h_{\phi}$ that maps embeddings to outputs: $g_{\theta}(X)\in\mathbb{R}^{d},\;h_{\phi}(g_{\theta}(X))=(\hat{y},\hat{s})\in\mathbb{R}^{2}$ (see Fig. 1). Two transfer learning strategies are considered in this study:
In Strategy 1 (Frozen encoder), the encoder $g_{\theta}$ is fixed and only the head $h_{\phi}$ is trained on paediatric data, with optimised parameters
$\phi^{*}=\arg\min_{\phi}\;\frac{1}{N_{C}}\sum_{j=1}^{N_{C}}\Big[\mathcal{L}_{\text{BCE}}(y_{j}^{C},\hat{y}_{j})+\mathcal{L}_{\text{MSE}}(s_{j}^{C},\hat{s}_{j})\Big],$ (4)
where $(\hat{y}_{j},\hat{s}_{j})=h_{\phi}(g_{\theta}(X_{j}^{C}))$ and $\mathcal{L}_{\text{BCE}}$ and $\mathcal{L}_{\text{MSE}}$ denote binary cross-entropy and mean squared error losses, respectively.
In Strategy 2 (Full fine-tuning), both encoder and head are updated:
$(\theta^{*},\phi^{*})=\arg\min_{\theta,\phi}\;\frac{1}{N_{C}}\sum_{j=1}^{N_{C}}\Big[\mathcal{L}_{\text{BCE}}(y_{j}^{C},\hat{y}_{j})+\mathcal{L}_{\text{MSE}}(s_{j}^{C},\hat{s}_{j})\Big],$ (5)
with $\theta^{*}$ and $\phi^{*}$ denoting the encoder and head parameters after fine-tuning on the paediatric dataset.
Table 1: Participant demographics and comparison of model performance (MAE and RMSE) across patients under various conditions. SNR reflects signal quality, where higher values indicate clearer snore signals. Scored AHI is based on home sleep tests scored by qualified sleep technologists. AHI Prediction: S/M: Single-/Multi-Task Learning; NF/F/UF: Non Fine-Tuning/Freezing/Un-Freezing Encoder; FD/SD: Fixed Delay (global median time delay)/Specific Delay (night-specific time delay).
ID Sex Age BMI Dur SNR Scored S- S- S- M- M- M- M-F M-UF M-F M-UF
(hrs) (dB) AHI NF F UF NF F UF -FD -FD -SD -SD
003 F 5 15.1 9.3 1.2 4 0.11 11.29 9.78 1.18 0.11 0.32 0.11 0.11 0.11 0.75
004 M 10 31.4 8.2 7.7 3.4 0.12 9.63 14.27 6.95 0.12 1.83 0.12 0.12 0.12 7.44
005 M 5 15.9 8.7 -3.1 9.4 0.11 1.49 5.17 3.79 0.34 0.57 0.11 0.11 0.11 1.03
006 F 2 17.8 10 5.5 0.8 0.6 9.5 4.1 0.3 0.3 9.7 0.3 0.1 5.5 0.7
007 M 11 33.1 7.8 1.5 6.4 0.13 1.15 0.64 2.18 0.64 6.41 2.31 0.13 6.41 0.13
008 F 9 22.99 7.7 -0.5 1.8 0.13 0.65 0.65 1.43 0.13 0.65 0.13 0.13 0.13 0.13
009 M 1 19.44 9.8 -6.8 8 0.1 1.53 1.53 0.51 0.71 2.55 0.1 0.1 0.1 0.1
010 M 1 16.77 8.8 -0.4 1.6 0.23 0.11 0.34 0.11 1.48 0.34 0.34 0.11 0.11 0.11
011 F 13 47.28 6.1 21 1.8 0.16 1.8 3.44 0.49 4.75 9.34 2.46 0.16 8.03 1.8
012 M 1 18.62 6.9 -0.4 2.6 0.14 1.3 0.43 1.88 1.3 0.14 0.43 3.04 0.72 0.14
013 M 5 17.08 8.3 1.4 5.3 0.12 0.36 2.53 0.12 1.33 1.57 3.25 4.46 2.53 4.46
014 M 8 28.45 9.6 -4.8 21 0.1 1.56 0.31 0.1 19.9 14.48 20.52 30.1 18.23 23.85
015 M 4 18.44 9.1 -18.9 0.4 0.33 1.43 0.77 0.33 0.55 0.77 0.33 0.11 0.11 0.11
016 F 15 41.61 5 13.0 0.8 0.2 1 16.2 1 0.6 0.2 8.2 0.2 3.8 1.4
017 M 6 32.15 8.6 -1.3 2.1 0.12 0.35 0.35 2.21 0.35 0.35 0.12 0.12 0.12 0.12
MAE 4.45 4.88 5.57 3.64 2.87 3.59 3.11 3.29 3.41 2.81
RMSE 6.81 6.89 7.92 6.3 3.9 4.68 4.2 4.55 4.26 3.86
4 EXPERIMENTAL SETUP
We first established two baselines: a single-task learning (STL) model and a multi-task learning (MTL) model trained on the adult dataset without fine-tuning, directly mapping acoustic features to OSA labels. The pretrained model jointly predicts OSA probability and SpO2 desaturation; in STL, the regression branch is omitted (Fig. 1).
Two fine-tuning strategies were then applied to the pretrained model (Section 3.4), yielding four variants across STL and MTL formulations. Within the MTL framework, we further incorporated SpO2 desaturation time-delay information during fine-tuning (Section 3.3). Two settings were considered: (1) a global median delay computed across all 15 paediatric subjects, and (2) subject-specific median delays computed per night. These produced four additional variants. In total, ten model configurations were evaluated.
4.1 Training setup
Adult models were trained for 50 epochs with a batch size of 1024 and a learning rate of $1\times 10^{-3}$ using the Adam optimiser. For paediatric adaptation, 15 nights of children’s Mel spectrograms with dual labels were fine-tuned for 20 epochs (batch size 8, learning rate $1\times 10^{-5}$, early stopping 5 epochs), balancing gradual updates and gradient noise to avoid overfitting. Models were implemented in TensorFlow Keras and trained on an NVIDIA RTX 8000 GPU (48GB).
4.2 Evaluation framework
The standard clinical measure for assessing OSA is the apnoea-hypopnoea index (AHI). This is defined as the number of apnea and hypopnea events per hour of sleep:
$\text{AHI}=\frac{N_{\text{apnea}}+N_{\text{hypopnea}}}{T_{\text{sleep}}}$ (6)
where $N_{\text{apnea}}$ and $N_{\text{hypopnea}}$ denote the number of apnoea and hypopnoea events, and $T_{\text{sleep}}$ is the total sleep time in hours. Predicted AHI was derived by aggregating the model-predicted OSA events across all segments and normalising by total sleep time. In paediatric patients, OSA severity is classified based on AHI thresholds: <<5 as normal/mild, 5–10 as moderate, and >>10 as severe [20].
We used 5-fold cross-validation on the 15-night paediatric dataset, using 12 nights for fine-tuning and 3 for testing in each fold. This ensures that training and test data come from different patients, providing cross-patient validation. Model performance was assessed using mean absolute error (MAE) and root mean square error (RMSE) between the predicted AHI and the reference AHI from the scored HSAT recordings.
5 RESULTS AND DISCUSSION
Table 1 summarises AHI prediction across model variants. Fine-tuning consistently improves performance. Without adaptation, adult STL model collapses to near-zero predictions, classifying all patients as normal and failing to capture OSA severity. In contrast, fine-tuned models yield substantially higher AHI values and better discrimination. Multi-task learning generally outperforms single-task learning across all settings, regardless of encoder freezing. This highlights the benefit of jointly modelling OSA classification and SpO2 regression.
Incorporating delayed SpO2 further reduces prediction error, with night-specific delays providing the largest gains. Among models without night-specific delays, freezing the encoder performs on par with, or better than, unfreezing, suggesting that retaining pretrained acoustic representations is advantageous when physiological alignment is limited. The best results are achieved by combining MTL with full fine-tuning and subject-specific delays (M–UF–SD), which yields the lowest MAE (2.81) and RMSE (3.86).
From a clinical perspective these improvements are meaningful. Severe OSA (AHI $\geq$ 10) warrants further evaluation, and the patient may be advised to undergo additional diagnostic procedures at a hospital. For example, for patient ID 014 (scored AHI = 21), MTL with fine-tuning correctly identified the case as severe, whereas STL and non-fine-tuned models misclassified it as normal/mild. The M–UF–SD variant (multi-task learning with unfreezing encoder and night-specific delay) demonstrates particular promise in clinical relevance. It accurately identifies normal/mild patients (AHI: 0.4–4; Prediction: 0.11–1.8), reducing false positives, and provides close estimates for moderate cases (e.g., ID 013, scored AHI 5.3 vs predicted AHI 4.46). The small difference (MAE and RMSE) indicates the model’s reliability in estimating OSA severity. While some moderate cases were underestimated (e.g., IDs 005 and 007), overall trends indicate that the model reliably captures OSA severity and could support triage decisions in paediatric screening.
A key limitation of this study is the small size of the paediatric dataset (15 patients), which constrains statistical power and may limit generalisability across broader populations. Variability in age, BMI, and signal quality further amplifies this challenge. However, this limitation directly motivates our approach: by leveraging large adult datasets for pretraining and then carefully adapting to paediatric cases, we demonstrate how transfer learning and physiological alignment can mitigate the scarcity of labelled paediatric sleep data. This strategy provides a scalable path toward developing robust paediatric OSA models in data-constrained clinical settings.
6 Conclusion
In this paper, we propose a transfer learning approach to improve the paediatric diagnosis of OSA severity by exploiting knowledge from adult apnoea detection and fine-tuning it on limited paediatric data. We systematically investigated both single- and multi-task learning frameworks, incorporating oxygen desaturation as a secondary task and considering potential time-delay cues. Experimental results demonstrate that the proposed physiology-informed transfer learning framework enhances diagnostic accuracy, supporting the feasibility of non-invasive, child-friendly smartphone-based screening for clinicians and patients. While the size of available paediatric datasets remains a limiting factor, this study illustrates how transfer learning can help mitigate data scarcity and provide a scalable route for developing clinically relevant paediatric models. Future research will extend this approach by incorporating additional low-cost physiological signals, such as accelerometer or respiratory effort data (thorax and IMU motion sensors illustrated in Fig. 1), and by validating the framework in larger paediatric cohorts to strengthen clinical translation. Taken together, this work highlights transfer learning as a practical bridge between abundant adult datasets and the pressing need for accurate, accessible diagnostic tools in paediatric sleep medicine.